You invested $7000 in two accounts paying 7% and 8% annual interest, respectively. If the total interest earned for the year was $537, how much was invested at each rate?
Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
7000
Interest----- 537.00

Part I 7.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 7000 -------------1
Interest on both investments
7.00% x + 8.00% y= 537
Multiply by 100
7 x + 8 y= 53700.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -49000.00
Add this to (2)
0 x 1 y= 4700
divide by 1
y = 4700
Part I 7.00% $ 2300
Part II 8.00% $ 4700

CHECK
2300 --------- 7.00% ------- 161.00
4700 ------------- 8.00% ------- 376.00
Total -------------------- 537.00

m.ananth@hotmail.ca